# Closed-loop Long-horizon Robotic Planning via Equilibrium Sequence Modeling

### v1

@@ 1 Introduction
This work proposes an simple learning framework for planning via self-refinement. Specifically, we formulate the self-refining process as a fixed-point problem that recursively refines the plan until the equilibrium point, as illustrated in Fig. 1. While the forward process of this fixed-point problem could be solved efficiently using root-finding methods, more interestingly, its backpropagation can be skipped since its gradient is explicated by the implicit function theorem (Krantz & Parks, 2002) as in deep equilibrium models (Bai et al., 2019; Geng & Kolter, 2023). Such an analytical technique allows end-to-end supervised training of self-refining LLM planners without the need for additional verifiers or reward models in reinforcement learning, which substantially enhances its simplicity and practicality. And after training, our equilibrium planner is capable of dynamically allocating more inference computation in the equilibrium solving process to achieve better planning performance.
@@
  • We present equilibrium sequence modeling a simple training approach for self-refining LLMs based on equilibrium models, allowing for end-to-end learning in a supervised manner.
@@ 2 Related Work
Deep Equilibrium Models (Bai et al., 2019) are infinite-depth neural networks specified by fixed-point problems $x^{*}=f_{\theta}(x^{*})$, where $f_{\theta}$ is an equilibrium layer. While their inference can take infinite steps by the fixed-point iteration, their gradients are estimated using implicit differentiation (Krantz & Parks, 2002) without backpropagating through all layers, thus enabling memory-efficient training. They have been extensively applied to tasks such as visual understanding (Bai et al., 2020; 2022) and image generation (Pokle et al., 2022; Geng et al., 2023; Bai & Melas-Kyriazi, 2024). In this paper, we apply the fixed-point formulation of deep equilibrium models to the self-refinement process in LLM planners, allowing for simple supervised training of LLMs to refine themselves.
@@ 3.1 Preliminaries on Self-Refinement
Recently, Welleck et al. (2023); Shinn et al. (2023); Kim et al. (2023); Madaan et al. (2023) proposed self-refinement, which employs an LLM to iteratively refine the previous LLM output. This strategy naturally addresses the above limitations, since it introduces bidirectional token dependency and a dynamic error correction mechanism. Formally, let $x_{t}$ denote a draft plan and $c_{t}$ denote context (e.g. environmental feedback), then planning may be viewed as a self-refinement process as follows:
@@ 3.2 Self-Refinement as An Equilibrium Model
Its proof is provided in Appendix A. It is noteworthy that the inverse Jacobian term $A=(I-\frac{\partial f_{\theta}}{\partial x^{*}})^{-1}$ within the above gradient is difficult to compute exactly, for which existing work often approximates through the damped fixed-point unrolling or the Neumann series (Geng et al., 2021b). For simplicity, we follow Fung et al. (2022); Geng et al. (2021a) to drop the inverse Jacobian term using $A\approx I$ (which is justified by Theorem 3.1 in Fung et al. (2022) under appropriate assumptions):
@@ 3.3 Equilibrium Models with Feedback
planner $f$, environment or world model Env, number of iterations $N$.
@@ 3.4 Practical Implementation
Internal Feedback from World Model. Due to inefficiency of interacting with the environment, we maintain a world model (Ha & Schmidhuber, 2018) to provide the necessary feedback in closed-loop planning. Our world model takes the environmental context, task instruction and current plan as inputs and predicts some basic types of feedback. This definition is slightly simpler than the commonly used world model, which requires simulation of the environmental states, and therefore may to easier to train. Concretely, we implement the world model with an LLM and finetune it on the planner’s equilibrium feedback over all iterations for better generalizability. And during inference, we alternate between using external and internal feedback in closed-loop robotic planning.
@@ 5 Conclusion
Appendix A Proof of Implicit Function Theorem

### v4

@@ 1 Introduction
This work proposes an simple learning framework for planning via self-refinement. Specifically, we formulate the self-refining process as a fixed-point problem that recursively refines the plan until the equilibrium point, as illustrated in Fig. 1. While the forward process of this fixed-point problem could be solved efficiently using root-finding methods, more interestingly, its backpropagation can be skipped since its gradient is explicated by the implicit function theorem (Krantz & Parks, 2002) as in deep equilibrium models (Bai et al., 2019; Geng & Kolter, 2023). It is noted that the derived gradient term may be further simplified through a Jacobian-free approximation (Fung et al., 2022) to facilitate training. These analytical techniques allow end-to-end supervised training of the LLM planner to accomplish self-refinement without the need for additional verifiers or reward models in reinforcement learning-based counterparts, greatly enhancing simplicity and practicality. And after training, our equilibrium model-based planner is capable of dynamically allocating more inference computation in the equilibrium solving process to achieve better planning performance.
@@
  • We present equilibrium sequence modeling, a simple training approach for self-refining LLMs based on equilibrium models, allowing for end-to-end learning in a supervised manner.
@@ 2 Related Work
Deep Equilibrium Models (Bai et al., 2019) are infinite-depth neural networks specified by fixed-point problems $x^{*}=f_{\theta}(x^{*})$, where $f_{\theta}$ is an equilibrium layer. While their inference can take infinite steps by the fixed-point iteration, their gradients are estimated using implicit differentiation (Krantz & Parks, 2002) without backpropagating through all layers, thus enabling memory-efficient training. They have been extensively applied to tasks such as visual understanding (Bai et al., 2020; 2022) and image generation (Pokle et al., 2022; Geng et al., 2023; Bai & Melas-Kyriazi, 2024). In this paper, we apply the fixed-point formulation of deep equilibrium models to the self-refinement process in LLM planners, allowing for simple supervised training of LLMs to refine themselves. More detailed introduction to deep equilibrium models is presented in Appendix A.
@@ 3.1 Preliminaries on Self-Refinement
Recently, Welleck et al. (2023); Shinn et al. (2023); Kim et al. (2023); Madaan et al. (2023) proposed self-refinement, which uses an LLM $f_{\theta}$ to iteratively refine the previous LLM output. This strategy naturally addresses the above limitations, since it introduces bidirectional token dependency and a dynamic error correction mechanism. Formally, let $x_{t}$ denote a draft plan and $c_{t}$ denote context (e.g. environmental feedback), then planning may be viewed as a self-refinement process as follows:
@@ 3.2 Self-Refinement as An Equilibrium Model
Its proof is given in Section A.4. It is noteworthy that the inverse Jacobian term $A=(I-\frac{\partial f_{\theta}}{\partial x^{*}})^{-1}$ within the above gradient is difficult to compute exactly, for which existing work often approximates through the damped fixed-point unrolling or the Neumann series (Geng et al., 2021b). For computa-tional efficiency, we drop the inverse Jacobian term using $A\approx I$ as in Fung et al. (2022); Geng et al. (2021a); Choe et al. (2023), the latter having been validated on Transformer-based LLMs:
@@ 3.3 Equilibrium Models with Feedback
planner $f_{\theta}$, environment or world model Env, number of iterations $N$.
@@ 3.4 Practical Implementation
Internal Feedback from World Model. Due to inefficiency of interacting with the environment, we construct a world model (Ha & Schmidhuber, 2018) to provide the necessary feedback in closed-loop planning. Our world model takes the environmental context, task instruction and current plan as inputs and predicts some basic types of feedback. This definition is slightly simpler than the commonly used world model, which requires simulation of the environmental states, and therefore may be easier to train. Concretely, we implement the world model with an LLM and finetune it on the planner’s equilibrium feedback over all iterations for better generalizability. And during inference, we alternate between using external and internal feedback in closed-loop robotic planning.
@@ 5 Conclusion
Appendix A Background
A.1 Deep Equilibrium Models
Traditional neural networks are constructed by explicitly stacking layers $f^{(i)}$, which can be limited in their expressiveness due to the fixed number of layers and predetermined forward process. Instead, implicit models are defined by an underlying dynamic system to be solved, such as an ordinary differential equation (Chen et al., 2018), a controlled differential equation (Kidger et al., 2020), a stochastic differential equation (Kidger et al., 2021) or a fixed-point problem (Bai et al., 2019).
Deep equilibrium models, first introduced in Bai et al. (2019), are a representative class of implicit models characterized by fixed-point problems. Given an input $c$ and a function $f_{\theta}(x,c)$ such as a Transformer block (Bai et al., 2019) or a Transformer (Geng et al., 2023), deep equilibrium models define infinite-level stacking of this function $x_{i+1}=f_{\theta}(x_{i},c)$ with $i=0,1,\ldots,L$ and $L\rightarrow\infty$ by solving the solution $x^{*}$ to the following fixed-point equation defined by $f_{\theta}$ and $c$:
$x^{*}=f_{\theta}(x^{*},c)$ (11)
The forward pass of deep equilibrium models is root solving for the fixed-point problem. A common choice is the fixed-point iteration method, which starts from an initial guess $x_{0}$ and iteratively applies the transformation $x_{t+1}=f_{\theta}(x_{t},c)$ until convergence. A sufficient condition for its convergence is if $f_{\theta}$ is a contraction mapping w.r.t. $x$, namely its Lipschitz constant is less than one (Banach, 1922), which could be relaxed by the well-posedness condition in El Ghaoui et al. (2021). More advanced root solvers include Broyden’s method (Broyden, 1965) or Anderson acceleration (Anderson, 1965).
A.2 Training Deep Equilibrium Models
Unlike traditional neural networks, whose gradient requires backpropagation through time (Werbos, 1990) at high memory and computational cost, the gradient of deep equilibrium models is computed analytically without differentiating over its forward pass. Given an equilibrium point $x^{*}=f_{\theta}(x^{*},c)$ and a loss function $L(x^{*},y)$, the loss gradient w.r.t. the model parameters $\theta$ is provided by the implicit function theorem (Krantz & Parks, 2002; Bai et al., 2019) as follows:
Its proof is given in Section A.4. Due to the challenge of exactly computing the inverse Jacobian term $A=(I-\frac{\partial f_{\theta}}{\partial x^{*}})^{-1}$ in the above gradient, existing work often approximate it via the damped fixed-point unrolling or the Neumann series (Geng et al., 2021b). Recently, Fung et al. (2022); Geng et al. (2021a) propose to drop the inverse Jacobian term using $A\approx I$, with the former justifying it in Theorem 3.1 with strong assumptions. In practice, this approximation has been used extensively in one-step gradient (Bolte et al., 2023; Luketina et al., 2016; Finn et al., 2017; Liu et al., 2019; Garima et al., 2020) and shown to be effective on Transformer-based LLMs (Choe et al., 2023).
A.3 Transformer-based Deep Equilibrium Models
Deep equilibrium models are initially proposed on Transformer architecture (Vaswani et al., 2017) for language modeling tasks (Bai et al., 2019). This seminal work considers a Transformer block as the basic unit $f_{\theta}$ in the equilibrium model. Then, Geng et al. (2021a) investigates improvements over the Transformer block by replacing self-attention with matrix decomposition. They also introduce one-step gradient based on the approximation of $A\approx I$ for efficiency and stability, assuming that the Lipschitz condition apply to a large number of matrix decomposition methods.
Recently, following the prevalence of Diffusion Transformers (Peebles & Xie, 2023), deep equilibrium models are extended to image generation tasks. Geng et al. (2023) propose generative equilibrium Transformers consisting of two modules, one using Transformer as the basic unit $f_{\theta}$ in the equilibrium model. Their method yields advanced one-step image generation results. (Bai & Melas-Kyriazi, 2024) replace most of the intermediate Transformer blocks with an equilibrium model, thus significantly reducing the number of parameters and memory usage for training and inference.
A.4 Proof of Implicit Function Theorem